OpenStudy (anonymous):

Find the solutions to the following quadratic equation using the quadratic formula: x^2 - 12x + 36 = 0 Write the answers as integers.

OpenStudy (help!!!!):

Yeah so do you know the quadratic formula?

OpenStudy (help!!!!):

@Keridavis

OpenStudy (anonymous):

x+x-12x+36=0

OpenStudy (anonymous):

then.. um.. i forgot what was next..

OpenStudy (anonymous):

Here u go .. just use this formula.,, http://www.oncalc.com/wp-content/uploads/2011/02/quadratic-equation-calculator.gif

OpenStudy (help!!!!):

so we know it is a quadratic because it is in this form, a^2+a+5

OpenStudy (anonymous):

so what a b and c? though/

OpenStudy (help!!!!):

sorry I'm confusing you, lets start over

OpenStudy (anonymous):

I'm so lost!

OpenStudy (help!!!!):

using x^2 - 12x + 36 = 0 ax2 + bx + c like so. Now use the quadratic formula

OpenStudy (anonymous):

\[\large ax^2+bc+c=0\\ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] it is a direct computation

OpenStudy (help!!!!):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

\[x^2 - 12x + 36 = 0 \\a=1,b=-12,c=36\]

OpenStudy (anonymous):

course it would be a lot easier to recognize \[x^2-12x+36=(x-6)^2\]and then set \[(x-6)^2=0\\ x-6=0\\x=6\]

OpenStudy (anonymous):

you will get \(x=6\) if you use the formula was well

OpenStudy (anonymous):

Thank you hun.

OpenStudy (anonymous):

yw dear